The pdf of a continuous random variable R is given by f(r)={kr for 0<x<4 both inclusive

{0,elsewhere

a)determine c hence compute p(10<r<2)both inclusive, E(r) and var(r)


1
Expert's answer
2021-07-16T13:31:43-0400

Solution:

(a): We know that "\int_a^bf(x)dx=1"

"\int_0^4(kr)dr=1"

"\Rightarrow [kr^2/2]_0^4=1
\\ \Rightarrow \dfrac12\times [16k]=1
\\ \Rightarrow 8k=1
\\ \Rightarrow k=\dfrac18"

(b):

So, "f(x)=\dfrac r8"

Now, "P(-10\le r\le2)=P(-10\le r<0)+P(0\le r\le 2)"

"=0+\int_0^2(\dfrac r8)dr
\\=\dfrac{r^2}{16}|_0^2
\\=\dfrac14"

Next, "E(r)=\int_0^4r(\dfrac r8)dr=\int_0^4(\dfrac {r^2}8)dr"

"=\dfrac{r^3}{24}|_0^4
\\=\dfrac{64}{24}-0
\\=\dfrac 83"

"E(r^2)=\int_0^4r^2(\dfrac r8)dr=\int_0^4(\dfrac {r^3}8)dr
\\ =\dfrac{r^4}{32}|_0^4
\\=\dfrac{16\times 16}{32}-0
\\=8"

Now, "Var(r)=E(r^2)-[E(r)]^2=8-(\dfrac83)^2=\dfrac89"